(1) If a disk has a mass of 200g and a radius of 11.3cm, what is the moment of inertia of the disc?

(2) What is the translational motion counterpart to the moment of inertia?
Answer #1

a)

Moment of inertia

I = 0.5* m r^2

I = 0.5* 0.2* 0.113^2

I = 1.277*10^-3 kgm^2

======

b)

mass is counterpart of moment of inertia in translation motion.

======